OpenStudy (anonymous):

Calculus 1: Show that the inequality is true.

OpenStudy (anonymous):

Show that\[0 \le \int\limits_{2}^{3}\sin \sqrt[3]{x}dx \le 1\]

OpenStudy (anonymous):

is this on a test

OpenStudy (anonymous):

Nope. Just homework.

OpenStudy (anonymous):

liar i had this already it is on a test